Please help me answer this question. I have been working on it for hours and am not even close to getting near any of the multiple choice answers. I found this site and it said this question was already answered, however, every time i click on the problem no answer shows up. [[[what is the detention in a sedimentation tank 80 feet long, 30 feet wide, and 10 feet deep, if the flow is 2.1MGD?]]] Thank you vey much,
Sherry

Answer by rothauserc(4718)

detention time = total volume / feed rate
:
2.1 MGD = 11697.0122 cu.ft./hr
:
detention time = (80 * 30 * 10) / 11697.0122 = 2.0518 hours
:
another way to do this is
:
(((80 * 30 * 10) * 7.48) / (2100000)) * 24 = 2.0517 hours
:
Note 1 cu ft = 7.48 gallons
:
Note MGD is millions of gallons per day
: